OpenStudy (anonymous):

help please find the 9th term of the sequence. a4=3/4 and a8=1/2 ; a9= ???

OpenStudy (anonymous):

Let a be the first term and d be c.d. tn=a+(n-1)d 3/4=a+(4-1)d=a+3d 1/2=a+(8-1)d=a+7d subtract to get the value of d and then a9=a8+d

OpenStudy (cwrw238):

i guess this is an arithmetic sequence?

OpenStudy (anonymous):

i did this problem twice already and i got two different answers. i got 3/4 and the second time i got 1/2

OpenStudy (anonymous):

i gave the the solution considering it as an arithmetic sequence. you can also solve it as any other sequence e.g.,geometric

OpenStudy (cwrw238):

is it arithmetic sequences you are studying istinkatmath1?

OpenStudy (anonymous):

yes @cwrw238

OpenStudy (anonymous):

\[4d=\frac{ 1 }{ 2 }-\frac{ 3 }{ 4 }=\frac{ 2-3 }{4 }=-\frac{ 1 }{ 4 },d=-\frac{ 1 }{ 16 }\] \[a9=\frac{ 1 }{ 2 }+\frac{ -1 }{ 16 }=\frac{ 8-1 }{ 16 }=\frac{ 7 }{ 16 }\]

OpenStudy (anonymous):

i got it thanks !(: